a choclate chip cookie recipe calls for 8 eggs to make 6 dozen cookies. In order to make 15 dozen cookies for a party, how many eggs will be needed?

Answer by stanbon(75887)

a choclate chip cookie recipe calls for 8 eggs to make 6 dozen cookies. In order to make 15 dozen cookies for a party, how many eggs will be needed?
-------
Use a proportion:
eggs/cookies = eggs/cookies
-----
x / 15doz = 8/6doz
----
x = 15(8/6)
-----
x = 120/6 = 20 eggs
-------------------------
Cheers,
Stan H.
========================